# The Race between Technology and Woman:

### v8

@@ 6.2.2 Contributions of factor inputs
The fifth to tenth columns report the changes attributable to each factor input. The expansion of ICT equipment contributes significantly to the increase in the income shares of male and female skilled labor and the decrease in the income shares of male and female unskilled labor. The magnitude of the effect depends on the extent to which the rate of increase in the wages of male or female skilled (unskilled) labor is greater (smaller) than that in the nominal output. A rise in male and female skilled labor also contributes modestly to changes in labor shares.
@@ 6.2.3 Effects of technological change
Table 6 presents the effects of three types of technological change on the relative wages and income shares of the four types of labor. The effects of technological change in Table 6 are measured in levels to facilitate interpretation of the magnitude of the effects on each relative wage and income share, whereas the contributions of factor inputs in Tables 4 and 5 are measured in logarithms to facilitate comparison of the contributions to the four wage gaps or four income shares. The first to third columns report the changes attributable to technological change embodied in ICT capital, technological change embodied in non-ICT capital, and sector-specific disembodied technological change, respectively, when labor is perfectly mobile across sectors. The effects of technological change embodied in ICT capital are significant, whereas those of technological change embodied in non-ICT capital and of sector-specific disembodied technological change are marginal. The direction of the effects of technological change embodied in ICT capital reflects the presence of capital–skill–gender complementarity and aligns with that of the effects of the expansion of ICT equipment in Tables 4 and 5. The results indicate that a fall in the relative price of ICT investment owing to technological progress contributes to a narrowing of the gender wage gap, a widening of the skill wage gap, an increase in the income share of skilled labor, and a decline in the income share of unskilled labor. The magnitude of the effect on labor shares is largest for male unskilled labor in percentage-point terms, while that of the effect on relative wages is largest for female skilled labor. The small effects of technological change embodied in non-ICT capital reflect the lack of significant decrease in the relative price of non-ICT investment and of complementarity among non-ICT capital, skills, and gender. The small effects of sector-specific disembodied technological change reflect the lack of significant change in sector-specific disembodied technology. The fourth to sixth columns report the changes attributable to the three types of technological change when labor is not perfectly mobile across sectors. The magnitude of the effects of technological change embodied in ICT capital remains largely unchanged regardless of whether labor is perfectly mobile across sectors.
@@
Perfect mobility Imperfect mobility
$\Delta\ln q_{i}$ $\Delta\ln q_{o}$ $\Delta\ln A_{n}$ $\Delta\ln q_{i}$ $\Delta\ln q_{o}$ $\Delta\ln A_{n}$
@@
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 891 –0. 015 0. 024 –0. 937 –0. 014 –0. 016
(0. 051) (0. 001) (0. 004) (0. 056) (0. 001) (0. 004)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 077 –0. 001 –0. 001 –0. 057 0. 002 –0. 093
(0. 006) (0. 000) (0. 002) (0. 007) (0. 000) (0. 001)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 264 0. 004 –0. 011 0. 235 0. 002 0. 090
(0. 014) (0. 000) (0. 002) (0. 015) (0. 000) (0. 001)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 289 0. 021 –0. 043 1. 339 0. 021 0. 005
(0. 052) (0. 001) (0. 004) (0. 057) (0. 001) (0. 004)
@@
$\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 006 0. 000 0. 000 0. 005 0. 000 0. 002
(0. 001) (0. 000) (0. 000) (0. 001) (0. 000) (0. 000)
$\Delta\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 017 0. 000 0. 000 0. 017 0. 000 0. 001
(0. 001) (0. 000) (0. 000) (0. 001) (0. 000) (0. 000)
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 039 0. 000 0. 002 –0. 036 0. 000 –0. 011
(0. 003) (0. 000) (0. 000) (0. 002) (0. 000) (0. 000)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 0. 000 0. 001 –0. 008 0. 000 0. 005
(0. 001) (0. 000) (0. 000) (0. 001) (0. 000) (0. 000)
Notes: The first to sixth columns report the changes attributable to technological change. The changes attributable to technological change in the first to third columns are computed using equations (21) and (22), while those in the fourth to sixth columns are computed using equations (25) and (26). Standard errors in parentheses are computed using bootstrap with 500 replications.
Table 7 presents the effects of technological change embodied in ICT capital on the relative wages and income shares of the four types of labor for each value of the labor supply elasticity. For ease of reference, the first column reproduces the first column of Table 6, which corresponds to the case in which the labor supply elasticity is zero. Given the difficulty of selecting a single value of the aggregate hours elasticity (Chetty, Guren, Manoli and Weber, 2011; Keane and Rogerson, 2012), the second to fourth columns show how the effects vary according to the labor supply elasticity in the range of 0.5 to 2. The effects of technological change embodied in ICT capital on the skilled gender wage gap and the female skill wage gap are greatest in absolute value when the supply of labor is exogenous and become smaller in absolute value as the supply of labor is more elastic. Progress in ICT raises the value of the marginal product of female skilled labor relative to other types of labor, and thus, the relative wages of female skilled labor. If the supply of labor is endogenous, the quantities of male (female) skilled labor relative to female skilled (unskilled) labor can decrease (increase) in response to progress in ICT. The decrease (increase) in the supply of male (female) skilled labor relative to female skilled (unskilled) labor moderates a fall (rise) in the skilled gender (female skill) wage gap. Such feedback effects become stronger as the supply of labor is more elastic. However, even if the labor supply elasticity is high at two, the effects of technological change embodied in ICT capital on the skilled gender wage gap and the female skill wage gap remain greater than their observed changes. Moreover, the effects of technological change embodied in ICT capital on the unskilled gender wage gap and the male skill wage gap do not vary substantially depending on the value of the labor supply elasticity. Even if the labor supply elasticity is high at two, the effects of technological change embodied in ICT capital on the unskilled gender wage gap and the male skill wage gap are greater than their observed changes. All else held constant, when the supply of labor is unit elastic, the fall in the relative price of ICT capital would have reduced the skilled (unskilled) gender wage gap by 49 (10) percentage points and raised the male (female) skill wage gap by 27 (77) percentage points.
Table 7: Effects of technological change embodied in ICT capital for different values of the elasticity of labor supply, 1980–2005
    $1/\gamma$ 0 1/2 1 2
Relative wages
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 891 –0. 633 –0. 490 –0. 336
(0. 051) (0. 043) (0. 036) (0. 028)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 077 –0. 093 –0. 100 –0. 103
(0. 006) (0. 009) (0. 010) (0. 013)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 264 0. 276 0. 269 0. 245
(0. 014) (0. 013) (0. 014) (0. 016)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 289 0. 963 0. 771 0. 553
(0. 052) (0. 045) (0. 039) (0. 032)
Labor shares
$\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 006 0. 010 0. 014 0. 020
(0. 001) (0. 001) (0. 001) (0. 002)
$\Delta\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 017 0. 019 0. 020 0. 022
(0. 001) (0. 001) (0. 001) (0. 001)
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 039 –0. 057 –0. 071 –0. 093
(0. 003) (0. 004) (0. 005) (0. 008)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 –0. 007 –0. 006 –0. 003
(0. 001) (0. 001) (0. 001) (0. 002)
@@
Notes: The changes in the levels of relative wages and labor shares attributable to technological change are computed using equations (23) and (24) for each value of the labor supply elasticity $1/\gamma$. Standard errors in parentheses are computed using bootstrap with 500 replications.
@@
The effects of technological change embodied in ICT capital on the income shares of the four types of labor tend to be smallest in absolute value when the supply of labor is exogenous and become greater in absolute value as the supply of labor is more elastic. The rate of change in the income share of labor of type $f$ can be decomposed as $\Delta\ln\Lambda_{f}=\Delta\ln(w_{f}/p)+\Delta\ln\ell_{f}-\Delta\ln y$. As the labor supply elasticity increases, the rate of increase in equilibrium wages becomes smaller, whereas the rates of increases in equilibrium quantities of labor and output become greater. The effects of progress in ICT on labor shares become greater (smaller) as the labor supply elasticity increases if the rate of increase in the equilibrium quantities of labor is large (small) relative to the rate of decrease in the equilibrium wages plus the rate of increase in the equilibrium quantities of output. The effects on the income shares of male and female skilled labor increase due to a greater increase in the equilibrium quantities of labor, whereas that on the income shares of male unskilled labor decrease due to a smaller increase in the equilibrium quantities of labor. The effect on the income share of female unskilled labor is insensitive to the value of the labor supply elasticity. All else held constant, the fall in the relative price of ICT capital would have raised the income share of male (female) skilled labor by 1.4 (2.0) percentage points and reduced the income share of male (female) unskilled labor by 7.1 (0.6) percentage points in the case in which the supply of labor is unit elastic. In this case, the fall in the relative price of ICT capital accounts for 34 (54) percent of the increase in the income share of male (female) skilled labor and 54 (31) percent of the decline in the income share of male (female) unskilled labor.
Table 6.2.3 presents the effects of technological change embodied in ICT capital on the relative wages and income shares of the four types of labor by country for the cases in which the supply of labor is perfectly inelastic and unit elastic. For ease of reference, the first column reproduces the first and third columns of Table 7. The second to last columns report the effects evaluated at the sample means over time for each country. The results indicate that progress in ICT narrows the gender wage gap and widens the skill wage gap for all countries, while it increases the income share of skilled labor and decreases the income share of unskilled labor. The magnitude of the effects of progress in ICT varies across countries, depending on the rate of decrease in the relative price of ICT investment, as seen in the bottom row of Table 6.2.3, and on the income share of ICT capital relative to the income shares of other factors, as described in Appendix A.2.5.
{sidewaystable}
[ph] Effects of technological change embodied in ICT capital for each country
All AUS AUT CZE DEU DNK FIN GBR ITA JPN NLD USA
Relative wages, $01/\gamma=$ 0
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 891 –1. 361 –0. 568 –0. 774 –0. 884 –3. 349 –0. 266 –0. 944 –0. 556 –1. 325 –1. 310 –0. 776
(0. 051) (0. 081) (0. 033) (0. 048) (0. 053) (0. 237) (0. 013) (0. 056) (0. 036) (0. 099) (0. 091) (0. 042)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 077 –0. 145 –0. 039 –0. 072 –0. 061 –0. 212 –0. 026 –0. 088 –0. 045 –0. 092 –0. 084 –0. 084
(0. 006) (0. 010) (0. 003) (0. 007) (0. 005) (0. 018) (0. 003) (0. 008) (0. 004) (0. 009) (0. 007) (0. 007)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 264 0. 417 0. 227 0. 310 0. 218 1. 113 0. 103 0. 227 0. 225 0. 191 0. 423 0. 178
(0. 014) (0. 022) (0. 014) (0. 016) (0. 011) (0. 078) (0. 005) (0. 012) (0. 016) (0. 008) (0. 023) (0. 008)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 289 1. 832 1. 041 1. 365 1. 245 5. 337 0. 305 1. 454 0. 834 1. 674 2. 253 0. 939
(0. 052) (0. 075) (0. 040) (0. 061) (0. 056) (0. 272) (0. 009) (0. 063) (0. 029) (0. 108) (0. 119) (0. 037)
Relative wages, $11/\gamma=$ 1
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 490 –0. 744 –0. 317 –0. 423 –0. 473 –1. 612 –0. 157 –0. 507 –0. 301 –0. 664 –0. 660 –0. 435
(0. 036) (0. 057) (0. 024) (0. 033) (0. 036) (0. 136) (0. 010) (0. 039) (0. 025) (0. 057) (0. 056) (0. 031)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 100 –0. 186 –0. 052 –0. 089 –0. 077 –0. 250 –0. 034 –0. 114 –0. 055 –0. 106 –0. 104 –0. 113
(0. 010) (0. 016) (0. 005) (0. 010) (0. 008) (0. 027) (0. 004) (0. 012) (0. 006) (0. 013) (0. 010) (0. 012)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 269 0. 432 0. 234 0. 310 0. 216 1. 007 0. 107 0. 233 0. 209 0. 181 0. 407 0. 193
(0. 014) (0. 022) (0. 014) (0. 018) (0. 012) (0. 068) (0. 006) (0. 012) (0. 012) (0. 011) (0. 022) (0. 010)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 0. 771 1. 099 0. 638 0. 812 0. 721 2. 870 0. 197 0. 849 0. 510 0. 882 1. 234 0. 571
(0. 039) (0. 056) (0. 032) (0. 045) (0. 040) (0. 166) (0. 008) (0. 045) (0. 023) (0. 062) (0. 075) (0. 028)
Labor shares, $01/\gamma=$ 0
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 039 –0. 075 –0. 025 –0. 027 –0. 024 –0. 108 –0. 013 –0. 037 –0. 032 –0. 032 –0. 051 –0. 036
(0. 003) (0. 005) (0. 002) (0. 002) (0. 002) (0. 008) (0. 001) (0. 002) (0. 002) (0. 002) (0. 004) (0. 002)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 –0. 010 –0. 006 –0. 007 –0. 005 –0. 026 –0. 005 –0. 006 –0. 005 –0. 004 –0. 004 –0. 008
Labor shares, $11/\gamma=$ 1
$\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 014 0. 014 0. 010 0. 011 0. 011 0. 028 0. 005 0. 010 0. 008 0. 012 0. 019 0. 010
(0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 003) (0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 002)
(0. 001) (0. 001) (0. 001) (0. 002) (0. 001) (0. 001) (0. 001) (0. 001) (0. 001) (0. 000) (0. 000) (0. 001)
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 071 –0. 130 –0. 047 –0. 047 –0. 043 –0. 177 –0. 024 –0. 067 –0. 053 –0. 052 –0. 085 –0. 068
(0. 005) (0. 008) (0. 003) (0. 004) (0. 003) (0. 014) (0. 002) (0. 005) (0. 004) (0. 004) (0. 006) (0. 005)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 006 –0. 008 –0. 005 –0. 007 –0. 004 –0. 017 –0. 006 –0. 004 –0. 002 –0. 003 0. 001 –0. 009
Progress in ICT
Notes: Country names are abbreviated as follows: AUS, Australia; AUT, Austria; CZE, the Czech Republic; DEU, Germany; DNK, Denmark; FIN, Finland; GBR, the United Kingdom; ITA, Italy; JPN, Japan; NLD, the Netherlands; and USA, the United States. The changes in the levels of relative wages and labor shares attributable to technological change are computed using equations (23) and (24) for each value of the labor supply elasticity $1/\gamma$. Standard errors in parentheses are computed using bootstrap with 500 replications.

### v9

@@ 6.2.2 Contributions of factor inputs
The fifth to last columns report the changes attributable to each factor input. The expansion of ICCT equipment contributes significantly to the increase in the income shares of male and female skilled labor and the decrease in the income shares of male and female unskilled labor. The magnitude of the effect depends on the extent to which the rate of increase in the wages of male or female skilled (unskilled) labor is greater (smaller) than that in the nominal output. A rise in male and female skilled labor also contributes modestly to changes in labor shares.
@@ 6.2.3 Effects of technological change
The first three columns of Table 7 present the effects of three types of technological change on the relative wages and income shares of the four types of labor. The effects of technological change in Table 7 are measured in levels to facilitate interpretation of the magnitude of the effects on each relative wage and income share, whereas the contributions of factor inputs in Tables 5 and 6 are measured in logarithms to facilitate comparison of the contributions to the four wage gaps or four income shares. The first to third columns report the changes attributable to technological change embodied in ICCT capital, technological change embodied in non-ICCT capital, and sector-specific disembodied technological change, respectively. The effects of technological change embodied in ICCT capital are significant, whereas those of technological change embodied in non-ICCT capital and of sector-specific disembodied technological change are marginal. These results hold not only for the four wage gaps and the four labor shares but also for the aggregate gender wage gap and the aggregate skill wage gap. The direction of the effects of technological change embodied in ICCT capital aligns with that of the effects of the expansion of ICCT equipment in Tables 5 and 6, reflecting the presence of capital–skill–gender complementarity. The results indicate that a fall in the relative price of ICCT investment owing to technological progress contributes to a narrowing of the gender wage gap, a widening of the skill wage gap, an increase in the income share of skilled labor, and a decline in the income share of unskilled labor. The magnitude of the effect on relative wages is largest for female skilled labor in percentage-point terms, while that of the effect on labor shares is largest for male unskilled labor. The small effects of technological change embodied in non-ICCT capital reflect the lack of significant decrease in the relative price of non-ICCT investment and of complementarity among non-ICCT capital, skills, and gender. The small effects of sector-specific disembodied technological change reflect the lack of significant change in sector-specific disembodied technology.
@@
$\Delta\ln q_{i}$ $\Delta\ln q_{o}$ $\Delta\ln A_{n}$ $\Delta\ln q_{i}$
    $1/\gamma$ 0 0 0 1/2 1 2
@@
$\Delta\left(\frac{w_{mh}}{w_{fh}}\right)$ –0. 886 –0. 015 0. 048 –0. 628 –0. 486 –0. 333
(0. 052) (0. 001) (0. 003) (0. 042) (0. 036) (0. 027)
$\Delta\left(\frac{w_{mu}}{w_{fu}}\right)$ –0. 074 –0. 001 0. 003 –0. 090 –0. 096 –0. 098
(0. 006) (0. 000) (0. 001) (0. 008) (0. 010) (0. 012)
$\Delta\left(\frac{w_{mh}}{w_{mu}}\right)$ 0. 262 0. 004 –0. 015 0. 272 0. 264 0. 239
(0. 014) (0. 000) (0. 001) (0. 013) (0. 013) (0. 015)
$\Delta\left(\frac{w_{fh}}{w_{fu}}\right)$ 1. 283 0. 021 –0. 070 0. 958 0. 767 0. 550
(0. 053) (0. 001) (0. 003) (0. 044) (0. 039) (0. 031)
$\Delta\left(\frac{\overline{w}_{M}}{\overline{w}_{F}}\right)$ –0. 249 –0. 004 0. 013 –0. 226 –0. 210 –0. 186
(0. 014) (0. 000) (0. 001) (0. 016) (0. 016) (0. 017)
$\Delta\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)$ 0. 564 0. 009 –0. 031 0. 465 0. 401 0. 319
(0. 013) (0. 000) (0. 001) (0. 015) (0. 016) (0. 017)
@@
$\Delta\left(\frac{w_{mh}\ell_{mh}}{py}\right)$ 0. 006 0. 000 0. 000 0. 010 0. 014 0. 019
(0. 001) (0. 000) (0. 000) (0. 001) (0. 001) (0. 001)
$\Delta\left(\frac{w_{fh}\ell_{fh}}{py}\right)$ 0. 017 0. 000 –0. 001 0. 019 0. 020 0. 022
(0. 001) (0. 000) (0. 000) (0. 001) (0. 001) (0. 001)
$\Delta\left(\frac{w_{mu}\ell_{mu}}{py}\right)$ –0. 039 –0. 001 0. 002 –0. 056 –0. 070 –0. 091
(0. 002) (0. 000) (0. 000) (0. 003) (0. 005) (0. 007)
$\Delta\left(\frac{w_{fu}\ell_{fu}}{py}\right)$ –0. 008 0. 000 0. 001 –0. 008 –0. 007 –0. 004
(0. 001) (0. 000) (0. 000) (0. 001) (0. 001) (0. 002)
$\Delta\left(\frac{\sum_{f}w_{f}\ell_{f}}{py}\right)$ –0. 001 0. 000 0. 000 –0. 006 –0. 009 –0. 013
(0. 004) (0. 000) (0. 000) (0. 004) (0. 004) (0. 004)
@@
Notes: The changes in the levels of relative wages and labor shares attributable to technological change in the first to third columns are computed using equations (21) and (22) of Proposition 3, those in the fourth to last columns are computed using equations (A1) and (A2) of Proposition 4 for each value of the labor supply elasticity $1/\gamma$. Standard errors in parentheses are computed using bootstrap with 500 replications.
The last three columns of Table 7 presents the effects of technological change embodied in ICCT capital on the relative wages and income shares of the four types of labor for each value of the labor supply elasticity. For ease of reference, the first column reproduces the first column of Table 7, which corresponds to the case in which the labor supply elasticity is zero. Given the difficulty of selecting a single value of the aggregate hours elasticity (Chetty, Guren, Manoli and Weber, 2011; Keane and Rogerson, 2012), the second to fourth columns show how the effects vary according to the labor supply elasticity in the range of 0.5 to 2. The effects of technological change embodied in ICCT capital on the skilled gender wage gap and the female skill wage gap are greatest in absolute value when the supply of labor is exogenous and become smaller in absolute value as the supply of labor is more elastic. Progress in ICCT raises the value of the marginal product of female skilled labor relative to other types of labor, and thus, the relative wages of female skilled labor. If the supply of labor is endogenous, the quantities of male (female) skilled labor relative to female skilled (unskilled) labor can decrease (increase) in response to progress in ICCT. The decrease (increase) in the supply of male (female) skilled labor relative to female skilled (unskilled) labor moderates a fall (rise) in the skilled gender (female skill) wage gap. Such feedback effects become stronger as the supply of labor is more elastic. However, even if the labor supply elasticity is high at two, the effects of technological change embodied in ICCT capital on the skilled gender wage gap and the female skill wage gap remain greater than their observed changes. Moreover, the effects of technological change embodied in ICCT capital on the unskilled gender wage gap and the male skill wage gap do not vary substantially depending on the value of the labor supply elasticity. Even if the labor supply elasticity is high at two, the effects of technological change embodied in ICCT capital on the unskilled gender wage gap and the male skill wage gap are greater than their observed changes. All else held constant, when the supply of labor is unit elastic, the fall in the relative price of ICCT capital would have reduced the skilled (unskilled) gender wage gap by 49 (10) percentage points and raised the male (female) skill wage gap by 26 (77) percentage points.
@@
The effects of technological change embodied in ICCT capital on the income shares of the four types of labor tend to be smallest in absolute value when the supply of labor is exogenous and become greater in absolute value as the supply of labor is more elastic. The rate of change in the income share of labor of type $f$ can be decomposed as $\Delta\ln\Lambda_{f}=\Delta\ln(w_{f}/p)+\Delta\ln\ell_{f}-\Delta\ln y$. As the labor supply elasticity increases, the rate of increase in equilibrium wages becomes smaller, whereas the rates of increases in equilibrium quantities of labor and output become greater. The effects of progress in ICCT on labor shares become greater (smaller) as the labor supply elasticity increases if the rate of increase in the equilibrium quantities of labor is large (small) relative to the rate of decrease in the equilibrium wages plus the rate of increase in the equilibrium quantities of output. The effects on the income shares of male and female skilled labor increase due to a greater increase in the equilibrium quantities of labor, whereas that on the income shares of male unskilled labor decrease due to a smaller increase in the equilibrium quantities of labor. The effect on the income share of female unskilled labor is insensitive to the value of the labor supply elasticity. All else held constant, the fall in the relative price of ICCT capital would have raised the income share of male (female) skilled labor by 1.4 (2.0) percentage points and reduced the income share of male (female) unskilled labor by 7.0 (0.7) percentage points in the case in which the supply of labor is unit elastic. In this case, the fall in the relative price of ICCT capital accounts for 34 (53) percent of the increase in the income share of male (female) skilled labor and 53 (34) percent of the decline in the income share of male (female) unskilled labor. The effect of technological change embodied in ICCT capital also becomes more visible as the labor supply elasticity increases, although it is obscured when aggregated across different types of labor.
Appendix F presents three sets of additional results on the effects of technological change. First, the magnitude of the effects of technological change embodied in ICCT capital remains largely unchanged regardless of whether labor is perfectly mobile across sectors (Table A2). Second, for all countries, technological change embodied in ICCT capital narrows the gender wage gap and widens the skill wage gap, while it increases the income share of skilled labor and decreases the income share of unskilled labor (Table A3). Finally, the magnitude of the changes in relative wages and income shares attributable to ICCT capital may differ depending on whether they are obtained by estimating and aggregating sectoral production function parameters or by assuming and estimating an aggregate production function (Table A4).